4.11

4.11 An embankment content constructed at a 10% mo
Answer #1

ANSWER

From theory we know that water content will be:

w=\frac{W_{w}}{W_{s}}x100

So, W_{w} is the weight of water and W_{s} is the weight of solids.

W_{w}=W_{s}\left ( \frac{w}{100} \right )

W_{w}=\gamma _{d}V_{s}\left ( \frac{w_{2}-w_{1}}{100} \right )

Rate of compacted soil per hour is V_{s}=250ccy

W_{w}=\gamma _{d}V_{s}\left ( \frac{w_{2}-w_{1}}{100} \right )

W_{w}=(2850)(250)\left ( \frac{10-5}{100} \right )

W_{w}=35625lb/hr

W_{w}=\frac{35625}{8.33}gallons/hr

W_{w}=4276.71gallons/hr

\mathbf{W_{w}=4277gallons/hr}   ANS.

Therefore, the supplied water per hour to increase the water content is \mathbf{4277gallons/hr} .

Regards!!!